OpenStudy (anonymous):

Solve the formula for the specified variable. P = s1 + s2 + s3 for s3

OpenStudy (a_clan):

s3=P-(s1+s2)

OpenStudy (anonymous):

a. s3 = P - s1 - s2 b. s3 = s1 + P - s2 c. s3 = P + s1 + s2 d. s3 = s1 + s2 - P

jhonyy9 (jhonyy9):

yes s3 = P-s1-s2